lisa can shovel snow from her driveway in 60 minutes. bill can do the same in 55 minutes. how long would it take lisa and bill to shovel the driveway if they worked together?
Found 3 solutions by Alan3354, DominoEfx, richwmiller:
Answer by Alan3354(69443)

lisa can shovel snow from her driveway in 60 minutes. bill can do the same in 55 minutes. how long would it take lisa and bill to shovel the driveway if they worked together?
---------------
Lisa does 1/60 per minute
Bill does 1/55 per minute
------------
What do they do together?
How long does it take working together?

Answer by DominoEfx(1)
Answer by richwmiller(17219)

1/60+1/55=1/x
Working together should be about 30 minutes
x = 660/23
x = 28.696